OpenStudy (anonymous):

9q2(q4) Simplify please.

OpenStudy (anonymous):

9q6

OpenStudy (anonymous):

Law of exponents : x^m * x^n = x^m+n , when the bases are same then we can add the exponents. So 9 q^2 (q^4) = 9 q^2* q^4 = 9 q^2+4 = 9q^6 . I hope this helps :) Thanks Abhishek Sinha